OpenStudy (anonymous):

Tania is 15 years older than Robert. In five years, Tania will be twice as old as Robert is now. How old is Tania right now? Options are: a: 30 b: 35 c: 40 d: 45

OpenStudy (cj49):

b is the ans

OpenStudy (anonymous):

how?

OpenStudy (cj49):

if tania is 35 then 35-15=20(roberts age) after 5 yrs tania wud be 35+5=40 i.e twice as old as robert is rn

OpenStudy (anonymous):

Alright Thanks :)

OpenStudy (cj49):

let the age of robert be x rn tania=x+15 in 5 yrs tania= x+20 x+20=2x x=20 substitute x in the first equation 20+15=35

OpenStudy (cj49):

np c:

OpenStudy (anonymous):

Your answer is right

OpenStudy (cj49):

i'm glad its right....

OpenStudy (anonymous):

Me Too :)